A Bibliography of Publications in Theoretical Computer Science: 1975–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

08 February 2014 Version 1.08

Title word cross-reference

(log n)^{j} [14]. 5 [157]. β [166]. d [39]. k [182]. λ [62, 166, 4, 74, 178]. λ − β − η [44]. M_{2}(K) [141]. N [105]. st [45]. x, y [141]. xy [141]. yx [141].

* [20, 100].

-ary [39]. -calculi [74]. -calculus [62, 166, 44, 4, 178]. -dimensional [157]. -numbering [45]. -rational [105]. -recognisable [182].

1/* [100].

2*2 [134]. 2*2-matrix [134]. absolutely [11]. abstract [156, 74, 131]. acceptance [137]. acceptor [110]. addition [30, 157, 122, 1]. additions [42]. adjacency [73]. AFL [41]. AFLs [145]. algebra [104]. Algebraic [167, 189, 37, 77, 176, 13, 31, 169]. algorithm [166, 4, 103, 70, 125, 93, 11, 131]. Algorithms [15, 68, 84, 128, 134, 141]. Almost [182]. alphabet [61]. ambiguity [6]. Analysis [70, 64, 84]. application [31]. applications [131]. approach [36, 130]. Approximate [68]. arithmetic [180]. Arithmetical [162]. arrangements [85]. ary [39]. assertions [155]. associated [194]. asynchronous [95, 64]. automata [148, 121, 40, 191, 187, 64].

balanced [53]. beta [31]. between [99, 174, 21, 163, 63, 38, 181]. Bibliography [113]. bilinear [128, 134, 141]. biological [32]. blind [146]. Boolean [10, 2, 48]. bound [23, 42, 1, 65]. bounded [66]. bounded-right-context [66]. boundedness [122]. bounds [61, 142, 39]. brackets [196].

calculating [103]. calculi [74]. calculus [62, 175, 166, 44, 4, 31, 9, 178]. Call [9]. Call-by-name [9]. call-by-value [9]. cardinal [103]. Categories [81, 150]. cellular [32]. certain [96, 133, 194, 39]. Chain [184, 81, 190]. chain-complete [81]. chains [1]. Characterization [44, 112, 168, 54]. characterizations [189, 27]. characterize [170]. checking [93]. Church [175, 36]. Circuit [49]. class [10, 24]. classes [5, 12, 38]. closed [196]. closure [77, 130]. codes [52, 7]. codings [127]. combinational [10, 18]. Combinatorial [60, 7]. combinatory [171]. Commutation [170]. commutative [105, 186]. comparison [8]. Complete [57, 51, 117, 144, 81, 177, 102, 16, 89]. complexities [5, 181]. Complexity [91, 71, 2, 161, 104, 123, 75, 17, 162, 10, 163, 29, 12, 18, 85, 3, 160, 132, 179, 131]. Computability [33]. computation [116, 162, 128, 134, 141]. computational [123, 3, 132]. computations [104, 114, 111, 42]. Computing [45, 160]. concepts [33]. Concerning [66, 20]. Conditions [53]. cone [100]. conjecture [183]. connectivity [80, 78]. considered [107]. construction [147]. constructions [150]. context [47, 79, 6, 66, 72]. context-free [47, 6, 72]. continuous [82]. control [99]. convergence [119]. convex [140]. Correctness [36]. counter [17]. counterexample [183]. covering [47, 122, 149]. Cyclic [102]. cylinder [172].

D [100]. D/sub [100]. D0L [55, 136, 153, 67]. D0L-systems [55]. D=R [52]. Data [156, 132, 99]. Davis [75]. decidability [55, 79, 195]. Decidable [80]. decision [56, 69, 126, 87, 154]. decomposition [13, 15]. definition [99]. definitions [119]. degree [6, 65]. degrees [29]. delay [7]. densest [118]. dependencies [99]. depth [49]. derivation [25]. description [83]. descriptions [92]. designs [94]. Deterministic [28, 172, 69, 5, 148, 57, 79]. developmental [5]. diagonalization [163]. differentiation [3]. dimensional [157]. direct [93]. discrimination [166, 171]. disjoint [48]. distributions [53]. divisions [142]. do [82]. domains [108]. DP0L [120]. DPDA’s [83]. DT0L [186]. dual [181].

E0L [117, 127]. Economy [83]. EDT0L [21]. effect [111]. Effectively [108]. efficient [87]. Eilenberg [106]. Elementary [136, 65, 178]. embedding [12]. enough [155]. enough-or [155]. enriched [150]. equations [96]. equivalence [69, 55, 120, 136, 185, 125, 86, 93, 67, 18]. erasable [26]. ergodic [187]. ET0L [21, 143]. Euclidean [89]. evaluate [142]. evaluating [180]. evaluation [71]. events [154]. examples [101]. expected [125]. explicit [147]. exponential [54]. exponential-time [54]. expressions [180, 39]. extension [164, 169]. extensions [13].

facets [109]. factorization [70]. Families [196, 26, 13, 41]. Fast [173, 11]. fields [169]. file [94]. Finding [140, 152]. Finite [40, 69, 121, 191, 103, 98, 187, 139, 7]. Finiteness [59]. firing [135]. Fixed [150, 35, 119]. Fixed-point [150]. fixpoint [188]. flowchart [189, 168]. form [120, 82, 86]. Formal [192, 99]. formation [161]. Forms [117, 120, 44, 110, 86, 143]. formulae [147]. foundations [60, 22]. FP0L [127]. free [47, 79, 6, 63, 72]. full [145]. Fully [74]. function [153]. functions [62, 112, 10, 147, 119, 12, 48, 67, 76, 38, 39, 179]. fundamental [111].

gaps [163]. generable [24]. generalised [151]. generalized [188, 68]. generating [170, 153]. generation [167, 173]. given [108]. global [116, 62, 174]. good [23, 143]. grammars [90, 185, 47, 190, 93, 149, 66]. grammatical [54]. graph [189, 16, 73]. group [104, 78, 128].

half [152]. half-spaces [152]. having [7]. HD0L [195]. hemisphere [118]. hierarchies [71]. hierarchy [159, 162, 50, 51]. holds [171]. homomorphisms [136, 28]. hyper [13]. hyper-algebraic [13]. hypothesis [161].

identification [46]. II [193, 129, 60, 134]. III [141]. images [133]. Improved [142]. improvement [69]. inclusion [19]. induction [192, 193]. inductive [46, 130]. ineffable [59]. inference [92, 46]. infinite [27, 103]. Information [27, 60, 22, 99, 23]. Information-theoretic [27]. inherent [6]. Init [158]. integers [71]. interactions [5, 32]. interpolating [3]. interpretation [31]. Intersection [133, 140, 152]. Intuitionistic [177]. inverse [44]. Isomorphism [120]. isomorphisms [144]. isotropy [128]. iterated [97]. Iterating [97]. iterative [167, 37].

K-calculus [31]. knapsack [68].

L [86, 139, 113]. L* [139]. L-forms [86]. labelled [31]. lambda [175, 31, 9]. lambda-calculus [31, 9]. Language [158, 96, 33, 61, 13, 139, 107]. languages [172, 37, 137, 124, 129, 5, 21, 127, 25, 19, 90, 26, 17, 79, 24, 54, 190, 138, 6, 196, 41, 28, 72]. large [173]. lattice [82]. LCF [101, 107]. Least [35]. left [149]. left-recursive [149]. lemma [72]. lemmas [14]. length [1]. letter [61]. Lindenmayer [194, 28]. linear [10, 47, 133]. linearity [125]. lists [109]. LL [93]. local [116]. Log [115, 144, 152, 41]. log-tape [144]. logic [177, 39]. Lower [39, 42, 142, 1]. lowness [123]. LR [90].

machines [69, 114, 146, 115]. many [109]. mapping [128]. mappings [126, 128, 134, 141]. match [94]. mathematical [60, 22]. matrices [103, 98, 73, 195]. matrix [151, 2, 134]. McCreight [131]. McCreight-Meyer [131]. meaningful [168]. measures [71]. mechanized [161]. Mechanizing [192, 193, 58]. method [151]. Meyer [131]. Minimal [29, 135]. minimum [11]. models [74]. monoid [139]. monoids [52, 63]. monotone [147, 2, 42, 183, 179]. monotonic [38]. mu [59]. mu-ineffable [59]. multicounter [146]. multihead [191]. multiple [115]. multiplication [40, 169, 134]. multiplications [142]. multiplications/divisions [142].

n [152, 41, 20]. N-subsets [20]. name [9]. natural [46]. ND [64]. ND-automata [64]. nearly [179]. necessary [142]. net [80]. nets [56, 91]. networks [135, 2, 183]. Non [94, 148]. non-singular [148]. Non-uniform [94]. nondeterministic [17]. nonlinear [49]. nonsingular [195]. nonterminals [28]. normal [44]. note [172, 148, 88]. NP [102, 16, 89]. NP-complete [102, 16, 89]. number [180, 153, 42, 142]. numbering [45]. numbers [111].

O [152]. Omega [114, 82, 58, 79, 184]. Omega-computations [114]. omega-context-free [79]. omega-order [58]. One [121, 191, 146]. One-way [191, 146]. open [196]. operations [138]. operator [96, 12]. operators [158]. optimal [92, 128, 134, 141]. oracle [115]. order [87, 154, 58, 150]. order-enriched [150]. ordering [102].

P [82, 184]. paging [85]. paging-complexity [85]. pairs [167, 37, 29]. parallel [25, 36]. Parikh [195]. parsers [83]. partial [94]. partial-match [94]. partially [146]. PASCAL [101]. PDA’s [83]. periodic [182, 85]. periodicity [174]. permanent [160]. Petri [56, 80, 91]. point [150]. points [35, 119]. polyautomata [135]. polyhedra [140]. polynomial [159, 126, 14, 57, 8, 29, 177, 50, 51]. polynomial-space [177]. polynomial-time [50, 51]. polynomials [71, 15, 142, 3]. posets [81, 130]. Positive [43]. possessing [44]. pre [41]. pre-AFL [41]. prefix [52]. preserving [34]. primes [173]. primitives [170]. problem [56, 55, 136, 19, 30, 157, 82, 47, 118, 135, 156, 89, 6, 149]. problems [56, 80, 126, 68, 16, 57, 91, 60, 122, 195]. procedure [69, 87, 75, 154]. processing [61]. produced [100]. Product [100, 2]. Program [176]. programming [33, 107]. programs [36]. proof [101, 165, 163, 65]. properties [123, 170, 46, 73, 184]. property [7]. propositional [177]. pumping [72]. Pushdown [110, 69, 148]. Putnam [75].

quadratic [179]. quality [30]. question [106].

Rational [112, 165, 37, 87, 105, 100, 106, 67, 63, 43]. reachability [56, 80, 157]. real [111]. Realizing [48]. recognisable [182]. recognizable [20]. recognizing [73]. recurrent [133]. Recursive [155, 62, 27, 119, 12, 54, 149, 178]. reducibilities [8]. reducible [168]. reduction [151, 171, 95]. referring [183]. registers [180]. Regular [52, 75, 154]. Regularity [34]. Regularity-preserving [34]. related [56, 158, 41]. relating [15]. relation [181]. Relations [163, 112, 63, 34]. Relationship [174, 21]. Relationships [38]. relative [137]. Remarks [17, 146, 105]. repetitions [174]. representation [62, 124, 129, 187, 132]. required [180]. resolution [75]. result [79]. retracts [82]. retrieval [60, 22]. retrieval-mathematical [60]. return [181]. revisited [35]. right [66]. Rosser [175, 36].

S. [106]. salesman [89]. schemes [189, 168, 32]. Schnorr [88, 183]. Scott [82]. second [159]. semantically [168]. semantics [101, 33, 176]. semi [145]. semi-AFLs [145]. semigroups [103, 52, 98]. separatedness [88]. sequence [55, 120, 136, 153]. sequences [123, 106, 43]. sequential [112, 145, 76]. series [105, 133, 194]. sets [80, 182, 170, 144, 48, 51]. setting [188]. short [147]. sigma [20]. Simple [190, 19, 191, 70, 125]. simplified [16]. single [61]. singular [148]. size [49, 39]. slices [170]. solution [136, 82]. solutions [96]. Some [56, 123, 126, 16, 26, 46, 131, 68, 32, 91, 24, 195]. sorting [23, 11]. space [14, 99, 115, 41, 177, 181]. spaces [152]. specification [156]. specified [137]. squad [135]. stability [32]. stable [11]. step [159]. storage [60, 22, 11]. Strassen [65]. Strategies [193, 175]. strings [27]. strong [80, 46, 72]. strong-identification [46]. structural [192, 193]. structure [99]. structured [189]. structures [77]. study [96]. sub [100]. subclasses [24]. subexponential [29]. subpatterns [153]. subsequential [112]. subsets [20]. substitution [97]. Subword [5]. Surface [25]. Switching [179]. symbol [117]. symbolic [3]. symmetric [147]. synchronization [170, 135, 7]. syntactic [52, 139]. Syntactical [96]. synthesis [64]. system [192]. systems [167, 30, 163, 157, 186, 95, 22, 122, 194, 113, 195, 55].

table [78]. tape [61, 144]. tapes [115]. Tautology [151]. testing [151, 173]. their [156]. theorem [171, 13, 12, 164]. theorems [165, 20]. theoretic [189, 27]. theory [116, 188, 87, 23, 58, 131]. time [14, 57, 135, 8, 54, 152, 50, 51]. transductions [165, 37, 145]. transformation [185]. transformations [176]. transitive [77]. Translational [14]. traveling [89]. tree [25]. trees [124, 129, 25, 84, 185, 53]. Turing [114, 6]. turn [69]. Two [20, 165, 117, 145, 140]. two-symbol [117]. two-way [145]. type [58]. typed [4, 74, 178]. types [156].

undecidable [30]. unification [4, 58]. uniform [94, 130]. uniformly [26].

Valiant [69]. value [9]. variables [48]. variant [76]. Varieties [138, 128, 134, 141]. various [170, 5]. vector [30, 157, 122]. verification [173]. verify [78]. visit [121].

way [121, 146, 145, 191]. weak [171]. weighted [53]. where [139, 141]. which [142]. whose [179]. Winograd [164]. without [5]. words [174].

Z [67]. Z-rational [67]. Zeros [67].

References

Schonhage:1975:LBL

[1] A. Schonhage. A lower bound for the length of addition chains. Theoretical Computer Science, 1(1):1–12, June 1975. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paterson:1975:CMN

[2] M. S. Paterson. Complexity of monotone networks for Boolean matrix product. Theoretical Computer Science, 1(1): 13–20, June 1975. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Strassen:1975:CCS

[3] V. Strassen. The computational complexity of symbolic differentiation of interpolating polynomials. Theoretical Computer Science, 1(1):21–25, June 1975. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huet:1975:UAT

[4] G. P. Huet. A unification algorithm for typed λ-calculus. Theoretical Computer Science, 1(1):27–57, June 1975. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:1975:SCV

[5] A. Ehrenfeucht, K. P. Lee, and G. Rozenberg. Subword complexities of various classes of deterministic developmental languages without interactions. Theoretical Computer Science, 1(1):59–75, June 1975. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reedy:1975:TDI

[6] A. Reedy and W. J. Savitch. The Turing degree of the inherent ambiguity problem for context-free languages. Theoretical Computer Science, 1(1):77–91, June 1975. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Restivo:1975:CPC

[7] A. Restivo. A combinatorial property of codes having finite synchronization delay. Theoretical Computer Science, 1(2):

95–101, December 1975. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ladner:1975:CPT

[8] R. E. Ladner, N. A. Lynch, and A. L. Selman. A comparison of polynomial time reducibilities. Theoretical Computer Science, 1(2):103–123, December 1975. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Plotkin:1975:CNC

[9] G. D. Plotkin. Call-by-name, call-by-value and the lambda-calculus. Theoretical Computer Science, 1(2):125–159, December 1975. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harper:1975:CBF

[10] L. H. Harper, W. N. Hsieh, and J. E. Savage. A class of Boolean functions with linear combinational complexity. Theoretical Computer Science, 1(2):161– 133, December 1975. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Preparata:1975:FSS

[11] F. P. Preparata. A fast stable sorting algorithm with absolutely minimum storage. Theoretical Computer Science, 1 (2):185–190, December 1975. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moll:1976:OET

[12] R. Moll. An operator embedding theorem for complexity classes of recursive functions. Theoretical Computer Science, 1(3):193–198, February 1976. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Leeuwen:1976:DTH

[13] J. Leeuwen and D. Wood. A decomposition theorem for hyper-algebraic extensions of language families. Theoretical Computer Science, 1(3):199–214, February 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Book:1976:TLP

[14] R. V. Book. Translational lemmas, polynomial time, and (log n)^{j} — space. Theoretical Computer Science, 1(3):215– 216, February 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mignotte:1976:ARD

[15] M. Mignotte. Algorithms relating to the decomposition of polynomials. Theoretical Computer Science, 1(3):227– 235, February 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garey:1976:SSN

[16] M. R. Garey and D. S. Johnson. Some simplified NP-complete graph problems. Theoretical Computer Science, 1(3):237– 267, February 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Greibach:1976:RCN

[17] S. A. Greibach. Remarks on the complexity of nondeterministic counter languages. Theoretical Computer Science, 1(4):269–288, April 1976. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schnorr:1976:CCE

[18] C. P. Schnorr. The combinational complexity of equivalence. Theoretical Computer Science, 1(4):289–295, April 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Friedman:1976:IPS

[19] E. P. Friedman. The inclusion problem for simple languages. Theoretical Computer Science, 1(4):297–316, April 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karhumaki:1976:TTC

[20] J. Karhumaki. Two theorems concerning recognizable N-subsets of sigma *. Theoretical Computer Science, 1(4):317– 323, April 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:1976:RBE

[21] A. Ehrenfeucht, G. Rozenberg, and S. Skyum. A relationship between ET0L and EDT0L languages. Theoretical Computer Science, 1(4):325–330, April 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marek:1976:ISR

[22] W. Marek and Z. Pawlak. Information storage and retrieval systems: mathematical foundations. Theoretical Computer Science, 1(4):331–354, April 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fredman:1976:HGI

[23] M. L. Fredman. How good is the information theory bound in sorting?

Theoretical Computer Science, 1(4):355– 361, April 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meznik:1976:SSC

[24] I. Meznik. On some subclasses of the class of generable languages. Theoretical Computer Science, 2(1):1–7, June 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Engelfriet:1976:STL

[25] J. Engelfriet. Surface tree languages and parallel derivation trees. Theoretical Computer Science, 2(1):9–27, June 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ginsburg:1976:SUE

[26] S. Ginsburg, J. Goldstine, and S. Greibach. Some uniformly erasable families of languages. Theoretical Computer Science, 2(1):29–44, June 1976. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chaitin:1976:ITC

[27] G. J. Chaitin. Information-theoretic characterizations of recursive infinite strings. Theoretical Computer Science, 2(1):45–48, June 1976. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vitanyi:1976:DLL

[28] P. M. B. Vitanyi. Deterministic Lindenmayer languages, nonterminals and homomorphisms. Theoretical Computer Science, 2(1):49–71, June 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Machtey:1976:MPP

[29] M. Machtey. Minimal pairs of polynomial degrees with subexponential complexity. Theoretical Computer Science, 2(1):73–76, June 1976. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hack:1976:QPV

[30] M. Hack. The quality problem for vector addition systems is undecidable. Theoretical Computer Science, 2(1):77– 95, June 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Levy:1976:AIL

[31] J.-J. Levy. An algebraic interpretation of the lambda beta K-calculus; and an application of a labelled lambda-calculus. Theoretical Computer Science, 2(1):97– 114, June 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Herman:1976:SSB

[32] G. T. Herman and A. Walker. On the stability of some biological schemes with cellular interactions. Theoretical Computer Science, 2(1):115–130, June 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Egli:1976:CCP

[33] H. Egli and R. L. Constable. Computability concepts for programming language semantics. Theoretical Computer Science, 2(2):133–145, ???? 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seiferas:1976:RPR

[34] J. I. Seiferas and R. McNaughton. Regularity-preserving relations. Theoretical Computer Science, 2(2):147– 154, ???? 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeBakker:1976:LFP

[35] J. W. De Bakker. Least fixed points revisited. Theoretical Computer Science, 2(2):155–181, ???? 1976. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rosen:1976:CPP

[36] B. K. Rosen. Correctness of parallel programs: the Church–Rosser approach. Theoretical Computer Science, 2(2):183– 207, ???? 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boasson:1976:ALI

[37] L. Boasson. Algebraic languages, iterative pairs and rational transductions. Theoretical Computer Science, 2(2):209– 223, ???? 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Trakhtenbrot:1976:RBC

[38] M. B. Trakhtenbrot. Relationships between classes of monotonic functions. Theoretical Computer Science, 2(2):228– 247, ???? 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vilfan:1976:LBS

[39] B. Vilfan. Lower bounds for the size of expressions for certain functions in d-ary logic. Theoretical Computer Science, 2 (2):249–269, ???? 1976. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:1976:FAM

[40] O. H. Ibarra, S. K. Sahni, and C. E. Kim. Finite automata with multiplication. Theoretical Computer Science, 2 (3):271–294, September 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Springsteel:1976:PAL

[41] F. N. Springsteel. On the pre-AFL of (log n) space and related families of languages. Theoretical Computer Science, 2(3):295–304, September 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schnorr:1976:LBN

[42] C. P. Schnorr. A lower bound on the number of additions in monotone computations. Theoretical Computer Science, 2(3):305–315, September 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Soittola:1976:PRS

[43] M. Soittola. Positive rational sequences. Theoretical Computer Science, 2(3):317– 322, September 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dezani-Ciancaglini:1976:CNF

[44] M. Dezani-Ciancaglini. Characterization of normal forms possessing inverse in the λ − β − η-calculus. Theoretical Computer Science, 2(3):323–337, September 1976. CODEN TCSCDI.

ISSN 0304-3975 (print), 1879-2294 (electronic).

Even:1976:CN

[45] S. Even and R. E. Tarjan. Computing an st-numbering. Theoretical Computer Science, 2(3):339–344, September 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Minicozzi:1976:SNP

[46] E. Minicozzi. Some natural properties of strong-identification in inductive inference. Theoretical Computer Science, 2(3):345–360, September 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hunt:1976:CPL

[47] H. B. Hunt, D. J. Rosenkrantz, and T. G. Szymanski. The covering problem for linear context-free grammars. Theoretical Computer Science, 2(3):361– 382, September 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paul:1976:RBF

[48] W. J. Paul. Realizing Boolean functions on disjoint sets of variables. Theoretical Computer Science, 2(3):383–396, September 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paterson:1976:CSN

[49] M. S. Paterson and L. G. Valiant. Circuit size is nonlinear in depth. Theoretical Computer Science, 2(3):397–400, September 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stockmeyer:1976:PTH

[50] L. J. Stockmeyer. The polynomial-time hierarchy. Theoretical Computer Science, 3(1):1–22, October 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wrathall:1976:CSP

[51] C. Wrathall. Complete sets and the polynomial-time hierarchy. Theoretical Computer Science, 3(1):23–33, October 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lallement:1976:RSD

[52] G. Lallement. Regular semigroups with D=R as syntactic monoids of prefix codes. Theoretical Computer Science, 3 (1):35–49, October 1976. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hotz:1976:CBT

[53] G. Hotz. Conditions for balanced trees on weighted distributions. Theoretical Computer Science, 3(1):51–59, October 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Monien:1976:RGC

[54] B. Monien. A recursive and a grammatical characterization of the exponentialtime languages. Theoretical Computer Science, 3(1):61–74, October 1976. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Culik:1976:DSE

[55] K. Culik, II. On the decidability of the sequence equivalence problem for D0Lsystems. Theoretical Computer Science, 3(1):75–84, October 1976. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Araki:1976:SDP

[56] T. Araki and T. Kasami. Some decision problems related to the reachability problem for Petri nets. Theoretical Computer Science, 3(1):85–104, October 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jones:1976:CPD

[57] N. D. Jones and W. T. Laaser. Complete problems for deterministic polynomial time. Theoretical Computer Science, 3(1):105–117, October 1976. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jensen:1976:MOO

[58] D. C. Jensen and T. Pietrzykowski. Mechanizing omega-order type theory through unification. Theoretical Computer Science, 3(2):123–171, November 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Park:1976:FMI

[59] D. Park. Finiteness is mu-ineffable. Theoretical Computer Science, 3(2):173– 181, November 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lipski:1976:ISR

[60] W. Lipski, Jr. Information storage and retrieval-mathematical foundations II. combinatorial problems. Theoretical Computer Science, 3(2):183–211, November 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hartmanis:1976:TBS

[61] J. Hartmanis and L. Berman. On tape bounds for single letter alphabet language processing. Theoretical Computer Science, 3(2):213–224, November 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barendregt:1976:GRR

[62] H. Barendregt. A global representation of the recursive functions in the λcalculus. Theoretical Computer Science, 3(2):225–242, November 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schutzenberger:1976:RRB

[63] M. P. Schutzenberger. On the rational relations between free monoids. Theoretical Computer Science, 3(2):243–259, November 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Starke:1976:ASA

[64] P. H. Starke. Analysis and synthesis of asynchronous ND-automata. Theoretical Computer Science, 3(2):261–266, November 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schonhage:1976:EPS

[65] A. Schonhage. An elementary proof for Strassen’s degree bound. Theoretical Computer Science, 3(2):267–272, November 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Szymanski:1976:CBR

[66] T. G. Szymanski. Concerning boundedright-context grammars. Theoretical Computer Science, 3(3):273–282, December 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ruohonen:1976:ZZR

[67] K. Ruohonen. Zeros of Z-rational functions and D0L equivalence. Theoretical Computer Science, 3(3):283–292, December 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chandra:1976:AAS

[68] A. K. Chandra, D. S. Hirschberg, and C. K. Wong. Approximate algorithms for some generalized knapsack problems. Theoretical Computer Science, 3(3):293– 304, December 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beeri:1976:IVD

[69] C. Beeri. An improvement on Valiant’s decision procedure for equivalence of deterministic finite turn pushdown machines. Theoretical Computer Science, 3(3):305–320, December 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Knuth:1976:ASF

[70] D. E. Knuth and L. T. Pardo. Analysis of a simple factorization algorithm. Theoretical Computer Science, 3(3):321– 348, December 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lipton:1976:CMH

[71] R. J. Lipton and D. Dobkin. Complexity measures and hierarchies for the evaluation of integers and polynomials. Theoretical Computer Science, 3(3):349– 357, December 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wise:1976:SPL

[72] D. S. Wise. A strong pumping lemma for context-free languages. Theoretical Computer Science, 3(3):359–369, December 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rivest:1976:RGP

[73] R. L. Rivest and J. Vuillemin. On recognizing graph properties from adjacency matrices. Theoretical Computer Science, 3(3):371–384, December 1976. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Milner:1977:FAM

[74] R. Milner. Fully abstract models of typed λ-calculi. Theoretical Computer Science, 4(1):1–22, February 1977. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galil:1977:CRR

[75] Z. Galil. On the complexity of regular resolution and the Davis–Putnam procedure. Theoretical Computer Science, 4 (1):23–46, February 1977. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schutzenberger:1977:VSF

[76] H. P. Schutzenberger. On a variant of sequential functions. Theoretical Computer Science, 4(1):47–57, February 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lehmann:1977:AST

[77] D. J. Lehmann. Algebraic structures for transitive closure. Theoretical Computer Science, 4(1):59–76, February 1977. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vuillemin:1977:HVC

[78] J. Vuillemin. How to verify the connectivity of a group table. Theoretical Computer Science, 4(1):77–82, February 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Linna:1977:DRD

[79] M. Linna. A decidability result for deterministic omega-context-free languages. Theoretical Computer Science, 4(1):83– 98, February 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Araki:1977:DPS

[80] T. Araki and T. Kasami. Decidable problems on the strong connectivity of Petri net reachability sets. Theoretical Computer Science, 4(1):99–119, February 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Markowsky:1977:CCC

[81] G. Markowsky. Categories of chaincomplete posets. Theoretical Computer Science, 4(2):125–135, April 1977. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hosono:1977:RPO

[82] C. Hosono and M. Sato. The retracts in P omega do not form a continuous lattice — a solution to Scott’s problem. Theoretical Computer Science, 4 (2):137–142, April 1977. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Geller:1977:EDP

[83] M. M. Geller, H. B. Hunt, III, T. G. Szymanski, and J. D. Ullman. Economy of description by parsers, DPDA’s, and PDA’s. Theoretical Computer Science, 4(2):143–153, April 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Francon:1977:AAT

[84] J. Francon. On the analysis of algorithms for trees. Theoretical Computer Science, 4(2):155–169, April 1977. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stork:1977:PCP

[85] H.-G. Stork. On the paging-complexity of periodic arrangements. Theoretical Computer Science, 4(2):171–197, April 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maurer:1977:FEF

[86] H. Maurer, Th. Ottmann, and A. Salomaa. On the form equivalence of Lforms. Theoretical Computer Science, 4 (2):199–225, April 1977. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferrante:1977:EDP

[87] J. Ferrante and J. R. Geiser. An efficient decision procedure for the theory of rational order. Theoretical Computer Science, 4(2):227–233, April 1977. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lehmann:1977:NSS

[88] D. J. Lehmann. A note on Schnorr’s separatedness. Theoretical Computer Science, 4(2):235–??, April 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Papadimitriou:1977:ETS

[89] C. H. Papadimitriou. The Euclidean traveling salesman problem is NPcomplete. Theoretical Computer Science, 4(3):237–244, June 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Geller:1977:LKG

[90] M. M. Geller and M. A. Harrison. On LR(k) grammars and languages. Theoretical Computer Science, 4(3):245– 276, June 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jones:1977:CSP

[91] N. D. Jones, L. H. Landweber, and Y. E. Lien. Complexity of some problems in Petri nets. Theoretical Computer Science, 4(3):277–299, June 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Daley:1977:IOD

[92] R. Daley. On the inference of optimal descriptions. Theoretical Computer Science, 4(3):301–319, June 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Olshansky:1977:DAC

[93] T. Olshansky and A. Pnueli. A direct algorithm for checking equivalence of LL(k) grammars. Theoretical Computer Science, 4(3):321–349, June 1977. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Burkhard:1977:NUP

[94] W. A. Burkhard. Non-uniform partialmatch file designs. Theoretical Computer Science, 5(1):1–23, August 1977. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kwong:1977:RAS

[95] Y. S. Kwong. On reduction of asynchronous systems. Theoretical Computer Science, 5(1):25–50, August 1977. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chottin:1977:SSC

[96] L. Chottin. Syntactical study of certain language solutions of operator equations. Theoretical Computer Science, 5 (1):51–84, August 1977. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Engelfriet:1977:IIS

[97] J. Engelfriet. Iterating iterated substitution. Theoretical Computer Science, 5 (1):85–100, August 1977. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mandel:1977:FSM

[98] A. Mandel and I. Simon. On finite semigroups of matrices. Theoretical Computer Science, 5(2):101–111, October 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cremers:1977:FDD

[99] A. B. Cremers and T. N. Hibbard. On the formal definition of dependencies between the control and information structure of a data space. Theoretical Computer Science, 5(2):113–128, October 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Latteux:1977:PRC

[100] M. Latteux. Product in the rational cone produced by D/sub 1/*. Theoretical Computer Science, 5(2):129–134, October 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aiello:1977:PLS

[101] L. Aiello, M. Aiello, and R. W. Weyhrauch. PASCAL in LCF: semantics and examples of proof. Theoretical Computer Science, 5(2):135–177, October 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galil:1977:CON

[102] Z. Galil and N. Megiddo. Cyclic ordering is NP-complete. Theoretical Computer Science, 5(2):179–182, October 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jacob:1977:ACC

[103] G. Jacob. An algorithm for calculating the cardinal, finite or infinite, of the semigroups of matrices. Theoretical Computer Science, 5(2):183–204, October 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Atkinson:1977:CGA

[104] M. D. Atkinson. The complexity of group algebra computations. Theoretical Computer Science, 5(2):205–209, October 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karhumaki:1977:RCR

[105] J. Karhumaki. Remarks on commutative N-rational series. Theoretical Computer Science, 5(2):211–217, October 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reutenauer:1977:QER

[106] C. Reutenauer. On a question of S.

Eilenberg (rational sequences). Theoretical Computer Science, 5(2):219– ??, October 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Plotkin:1977:LCP

[107] G. D. Plotkin. LCF considered as a programming language. Theoretical Computer Science, 5(3):223–255, December 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Smyth:1977:EGD

[108] M. B. Smyth. Effectively given domains.

Theoretical Computer Science, 5(3):257– 274, December 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elgot:1977:MFL

[109] C. C. Elgot and L. Snyder. On the many facets of lists. Theoretical Computer Science, 5(3):275–305, December 1977. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ginsburg:1977:PAF

[110] S. Ginsburg and E. H. Spanier. Pushdown acceptor forms. Theoretical Computer Science, 5(3):307–320, December 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Luckhardt:1977:FEC

[111] H. Luckhardt. A fundamental effect in computations on real numbers. Theoretical Computer Science, 5(3):321–324, December 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Choffrut:1977:RRC

[112] C. Choffrut. Rational relations characterization of sequential and subsequential functions. Theoretical Computer Science, 5(3):325–337, December 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:1977:BS

[113] G. Rozenberg, M. Penttonen, and A. Salomaa. Bibliography of L systems. Theoretical Computer Science, 5(3):339–354, December 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cohen:1978:OCT

[114] R. S. Cohen and A. Y. Gold. Omegacomputations on Turing machines. Theoretical Computer Science, 6(1):1–23, February 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lynch:1978:LSM

[115] N. Lynch. Log space machines with multiple oracle tapes. Theoretical Computer Science, 6(1):25–39, February 1978. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abelson:1978:TTL

[116] H. Abelson. Towards a theory of local and global in computation. Theoretical Computer Science, 6(1):41–67, February 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Culik:1978:TSC

[117] K. Culik, II, H. A. Maurer, and Th.

Ottmann. On two-symbol complete E0L forms. Theoretical Computer Science, 6 (1):69–92, February 1978. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Johnson:1978:DHP

[118] D. S. Johnson and F. P. Preparata. The densest hemisphere problem. Theoretical Computer Science, 6(1):93–107, February 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manna:1978:CFF

[119] Z. Manna and A. Shamir. The convergence of functions to fixed points of recursive definitions. Theoretical Computer Science, 6(2):109–141, April 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Culik:1978:IFE

[120] K. Culik, II, H. A. Maurer, Th.

Ottmann, K. Ruohonen, and A. Salomaa. Isomorphism, form equivalence and sequence equivalence of DP0L forms. Theoretical Computer Science, 6 (2):143–173, April 1978. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Greibach:1978:OWF

[121] S. A. Greibach. One way finite visit automata. Theoretical Computer Science, 6(2):175–221, April 1978. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rackoff:1978:CBP

[122] C. Rackoff. The covering and boundedness problems for vector addition systems. Theoretical Computer Science, 6 (2):223–231, April 1978. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bennison:1978:SLP

[123] V. L. Bennison and R. I. Soare. Some lowness properties and computational complexity sequences. Theoretical Computer Science, 6(3):233–254, June 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Courcelle:1978:RTLa

[124] B. Courcelle. A representation of trees by languages. I. Theoretical Computer Science, 6(3):255–279, June 1978. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Knuth:1978:ELS

[125] D. E. Knuth and A. Schonhage. The expected linearity of a simple equivalence algorithm. Theoretical Computer Science, 6(3):281–315, June 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bollman:1978:SDP

[126] D. Bollman and M. Laplaza. Some decision problems for polynomial mappings. Theoretical Computer Science, 6 (3):317–325, June 1978. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:1978:ELC

[127] A. Ehrenfeucht and G. Rozenberg. E0L languages are not codings or FP0L languages. Theoretical Computer Science, 6(3):327–341, June 1978. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deGroote:1978:VOAa

[128] H. F. de Groote. On varieties of optimal algorithms for the computation of bilinear mappings. I. the isotropy group of a bilinear mapping. Theoretical Computer Science, 7(1):1–24, August 1978. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Courcelle:1978:RTLb

[129] B. Courcelle. A representation of trees by languages. II. Theoretical Computer Science, 7(1):25–55, August 1978. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wright:1978:UAI

[130] J. B. Wright, E. G. Wagner, and J. W.

Thatcher. A uniform approach to inductive posets and inductive closure. Theoretical Computer Science, 7(1):57– 77, August 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanEmdeBoas:1978:SAM

[131] P. van Emde Boas. Some applications of the McCreight-Meyer algorithm in abstract complexity theory. Theoretical Computer Science, 7(1):79–98, August 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Verbeek:1978:DRC

[132] R. Verbeek and K. Weihrauch. Data representation and computational complexity. Theoretical Computer Science, 7 (1):99–116, August 1978. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mignotte:1978:IIC

[133] M. Mignotte. Intersection of images of certain recurrent linear series. Theoretical Computer Science, 7(1):117– 121, August 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deGroote:1978:VOAb

[134] H. F. de Groote. On varieties of optimal algorithms for the computation of bilinear mappings. II. optimal algorithms for 2*2-matrix multiplication. Theoretical Computer Science, 7(2):127–148, October 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kobayashi:1978:MFT

[135] K. Kobayashi. On the minimal firing time of the firing squad synchronization problem for polyautomata networks. Theoretical Computer Science, 7(2):149–167, October 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrenfeucht:1978:EHS

[136] A. Ehrenfeucht and G. Rozenberg. Elementary homomorphisms and a solution of the D0L sequence equivalence problem. Theoretical Computer Science, 7(2):169–183, October 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Book:1978:LSR

[137] R. V. Book and C. Wrathall. On languages specified by relative acceptance. Theoretical Computer Science, 7(2):185– 195, October 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Perrot:1978:VLO

[138] J.-F. Perrot. Varieties of languages and operations. Theoretical Computer Science, 7(2):197–210, October 1978. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pin:1978:SMW

[139] J. E. Pin. On the syntactic monoid of L* where L is a finite language. Theoretical Computer Science, 7(2):211–215, October 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muller:1978:FIT

[140] D. E. Muller and F. P. Preparata. Finding the intersection of two convex polyhedra. Theoretical Computer Science, 7(2):217–236, October 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deGroote:1978:VOAc

[141] H. F. de Groote. On varieties of optimal algorithms for the computation of bilinear mappings. III. optimal algorithms for the computation of xy and yx where x, y in M_{2}(K). Theoretical Computer Science, 7(3):239–249, December 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schnorr:1978:ILB

[142] C. P. Schnorr. Improved lower bounds on the number of multiplications/ divisions which are necessary to evaluate polynomials. Theoretical Computer Science, 7(3):251–261, December 1978. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Skyum:1978:GEF

[143] S. Skyum. On good ET0L forms. Theoretical Computer Science, 7(3):263–272, December 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hartmanis:1978:LTI

[144] J. Hartmanis. On log-tape isomorphisms of complete sets. Theoretical Computer Science, 7(3):273–286, December 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:1978:TWS

[145] O. H. Ibarra. On two-way sequential transductions of full semi-AFLs. Theoretical Computer Science, 7(3):287–309, December 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Greibach:1978:RBP

[146] S. A. Greibach. Remarks on blind and partially blind one-way multicounter machines. Theoretical Computer Science, 7(3):311–324, December 1978. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kleiman:1978:ECS

[147] M. Kleiman and N. Pippenger. An explicit construction of short monotone formulae for the monotone symmetric functions. Theoretical Computer Science, 7(3):325–332, December 1978. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Friedman:1978:NNS

[148] E. P. Friedman. A note on non-singular deterministic pushdown automata. Theoretical Computer Science, 7(3):333– 339, December 1978. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Soisalon-Soininen:1979:CPL

[149] E. Soisalon-Soininen. On the covering problem for left-recursive grammars. Theoretical Computer Science, 8(1):1– 11, February 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wand:1979:FPC

[150] M. Wand. Fixed-point constructions in order-enriched categories. Theoretical Computer Science, 8(1):13–30, February 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bibel:1979:TTG

[151] W. Bibel. Tautology testing with a generalised matrix reduction method. Theoretical Computer Science, 8(1):31– 44, February 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Preparata:1979:FIH

[152] F. P. Preparata and D. E. Muller. Finding the intersection of n half-spaces in time O(n log n). Theoretical Computer Science, 8(1):45–55, February 1979. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Johansen:1979:GFN

[153] P. Johansen. The generating function of the number of subpatterns of a D0L sequence. Theoretical Computer Science, 8 (1):57–68, February 1979. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashiguchi:1979:DPO

[154] K. Hashiguchi. A decision procedure for the order of regular events. Theoretical Computer Science, 8(1):69–72, February 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Apt:1979:RAE

[155] K. R. Apt, J. A. Bergstra, and L. G. L. T. Meertens. Recursive assertions are not enough-or are they? Theoretical Computer Science, 8(1):73–87, February 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Majster:1979:DTA

[156] M. E. Majster. Data types, abstract data types and their specification problem. Theoretical Computer Science, 8 (1):89–127, February 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hopcroft:1979:RPD

[157] J. Hopcroft and J.-J. Pansiot. On the reachability problem for 5-dimensional vector addition systems. Theoretical Computer Science, 8(2):135–159, April 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Prodinger:1979:LOR

[158] H. Prodinger and F. J. Urbanek. Language operators related to init. Theoretical Computer Science, 8(2):161– 175, April 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baker:1979:SST

[159] T. P. Baker and A. L. Selman. A second step toward the polynomial hierarchy. Theoretical Computer Science, 8(2): 177–187, April 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Valiant:1979:CCP

[160] L. G. Valiant. The complexity of computing the permanent. Theoretical Computer Science, 8(2):189–201, April 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pudlak:1979:CMH

[161] P. Pudlak and F. N. Springsteel. Complexity in mechanized hypothesis formation. Theoretical Computer Science, 8 (2):203–225, April 1979. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hajek:1979:AHC

[162] P. Hajek. Arithmetical hierarchy and complexity of computation. Theoretical Computer Science, 8(2):227–237, April 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hartmanis:1979:RBD

[163] J. Hartmanis. Relations between diagonalization, proof systems, and complexity gaps. Theoretical Computer Science, 8(2):239–253, April 1979. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Morgenstern:1979:EWT

[164] J. Morgenstern. An extension of Winograd’s theorem. Theoretical Computer Science, 8(2):255–259, April 1979. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arnold:1979:NPT

[165] A. Arnold and M. Latteux. A new proof of two theorems about rational transductions. Theoretical Computer Science, 8(2):261–263, April 1979. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bohm:1979:DAI

[166] C. Bohm, M. Dezani-Ciancaglini, P. Peretti, and S. R. D. Rocca. A discrimination algorithm inside λ-β-calculus. Theoretical Computer Science, 8(3):271–291, June 1979. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beauquier:1979:AGS

[167] J. Beauquier. Algebraic generation and systems of iterative pairs. Theoretical Computer Science, 8(3):293–323, June 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elgot:1979:SMC

[168] C. C. Elgot and J. C. Shepherdson. A semantically meaningful characterization of reducible flowchart schemes. Theoretical Computer Science, 8(3):325– 357, June 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Winograd:1979:MAE

[169] S. Winograd. On multiplication in algebraic extension fields. Theoretical Computer Science, 8(3):359–377, June 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dolev:1979:CPG

[170] D. Dolev. Commutation properties and generating sets characterize slices of various synchronization primitives. Theoretical Computer Science, 8(3):379– 391, June 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hindley:1979:DTH

[171] R. Hindley. The discrimination theorem holds for combinatory weak reduction. Theoretical Computer Science, 8 (3):393–394, June 1979. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Autebert:1979:NCD

[172] J.-M. Autebert. A note on the cylinder of deterministic languages. Theoretical Computer Science, 8(3):395–399, June 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Plaisted:1979:FVT

[173] D. A. Plaisted. Fast verification, testing, and generation of large primes. Theoretical Computer Science, 9(1):1–16, July 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duval:1979:RBG

[174] J.-P. Duval. Relationship between global periodicity and repetitions of words. Theoretical Computer Science, 9(1):17– 26, July 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bergstra:1979:CRS

[175] J. Bergstra and J. W. Klop. Church– Rosser strategies in the lambda calculus. Theoretical Computer Science, 9(1): 27–38, July 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guessarian:1979:PTA

[176] I. Guessarian. Program transformations and algebraic semantics. Theoretical Computer Science, 9(1):39–65, July 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Statman:1979:IPL

[177] R. Statman. Intuitionistic propositional logic is polynomial-space complete. Theoretical Computer Science, 9 (1):67–72, July 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Statman:1979:TCE

[178] R. Statman. The typed λ-calculus is not elementary recursive. Theoretical Computer Science, 9(1):73–81, July 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wegener:1979:SFW

[179] I. Wegener. Switching functions whose monotone complexity is nearly quadratic. Theoretical Computer Science, 9(1):83–97, July 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flajolet:1979:NRR

[180] P. Flajolet, J. C. Raoult, and J. Vuillemin. The number of registers required for evaluating arithmetic expressions. Theoretical Computer Science, 9(1):99–125, July 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wechsung:1979:RBS

[181] G. Wechsung and A. Brandstadt. A relation between space, return and dual return complexities. Theoretical Computer Science, 9(1):127–140, July 1979. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Christol:1979:ARP

[182] G. Christol. Almost k-recognisable periodic sets. Theoretical Computer Science, 9(1):141–145, July 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wegener:1979:CCS

[183] I. Wegener. A counterexample to a conjecture of Schnorr referring to monotone networks. Theoretical Computer Science, 9(1):147–150, July 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tang:1979:CPP

[184] A. Tang. Chain properties in P omega.

Theoretical Computer Science, 9(2):153– 172, August 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harrison:1979:EGT

[185] M. A. Harrison, I. M. Havel, and A. Yehudai. On equivalence of grammars through transformation trees. Theoretical Computer Science, 9(2):173– 205, August 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Karhumaki:1979:CDS

[186] J. Karhumaki. On commutative DT0L systems. Theoretical Computer Science, 9(2):207–220, August 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Perrin:1979:ERF

[187] D. Perrin. The ergodic representation of finite automata. Theoretical Computer Science, 9(2):221–241, August 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ashcroft:1979:GSF

[188] E. A. Ashcroft and F. E. Fich. A generalized setting for fixpoint theory. Theoretical Computer Science, 9(2):243– 256, August 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bloom:1979:AGT

[189] S. L. Bloom and R. Tendell. Algebraic and graph theoretic characterizations of structured flowchart schemes. Theoretical Computer Science, 9(2):265– 286, October 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nijholt:1979:SCG

[190] A. Nijholt. Simple chain grammars and languages. Theoretical Computer Science, 9(2):287–309, October 1979. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Inoue:1979:OWS

[191] K. Inoue, I. Takanami, A. Nakamura, and T. Ae. One-way simple multihead finite automata. Theoretical Computer Science, 9(2):311–328, October 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aubin:1979:MSIa

[192] R. Aubin. Mechanizing structural induction. I. formal system. Theoretical Computer Science, 9(2):329–346, October 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aubin:1979:MSIb

[193] R. Aubin. Mechanizing structural induction. II. strategies. Theoretical Computer Science, 9(2):347–362, October 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reutenauer:1979:ASC

[194] C. Reutenauer. On the associated series to certain Lindenmayer systems. Theoretical Computer Science, 9(2):363– 375, October 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ruohonen:1979:SDP

[195] K. Ruohonen. On some decidability problems for HD0L systems with nonsingular Parikh matrices. Theoretical Computer Science, 9(2):377–384, October 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rodriguez:1979:FCL

[196] F. Rodriguez. Families of closed languages by open brackets. Theoretical Computer Science, 9(2):385–398, October 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).